OpenStudy (anonymous):

For each of the following, either use the subspace test to show that the given subset, W, is a subspace of V , or explain why the given subset is not a subspace of V . (b) V = lR^3 and W ={ (x, y, z) in V l x(y -z) = 0}.

OpenStudy (zugzwang):

It isn't a subspace. Here's a hint: Look for two elements of W (which would also be elements of IR³) Such that if you add them, the sum is not in W.

OpenStudy (zugzwang):

A stronger hint :D What does it mean, if x(y - z) = 0 This is the product of a pair of real numbers, so it can only mean that either x = 0 or (y - z) = 0 So, take both cases, so to speak... (If all that made sense to you :D )

OpenStudy (zugzwang):

All this says ONE If x = 0, it doesn't matter what y and z are, (x , y , z) ∈ W TWO If y = z, it doesn't matter what x is, (x , y , z) ∈ W So take an element that fits the first case, and take another element that fits the other...

OpenStudy (zugzwang):

Ok... (0 , 5 , 3) ∈ W Since 0(5 - 3) = 0 (2 , 2 , 2) ∈ W Since 2(2 - 2) = 0 But (0 , 5 , 3) + (2 , 2 , 2) = (2 , 7 , 5) 2(7 - 5) = 4 ≠ 0 Hence (0 , 5 , 3) + (2 , 2 , 2) = (2 , 7 , 5) ∉ W And therefore, W is not closed under vector addition; It cannot be a vector space; it cannot be a subspace of anything. //

OpenStudy (anonymous):

so a counter example would be say let v=(0,1,2), u=(1,1,1), thus u+v=(1,2,3) and hence x(y-z) not = 0. Would that be enough to explain its not a subspace?

OpenStudy (zugzwang):

Well, yes, it is. Something that is not a vector space in the first place, it cannot be a subspace :)

OpenStudy (anonymous):

haha yh you beat me to it. can you help me with the same question but about polynomials?

OpenStudy (zugzwang):

I'll do what I can.

OpenStudy (anonymous):

ok thanks let me type it.

OpenStudy (anonymous):

V=P_2(lR)(real polynomial functions of degree at most 2) and W={f inV lf(2)=2f(1)}

OpenStudy (anonymous):

W={f inV l f(2)=2f(1)} *

OpenStudy (zugzwang):

Notation's getting me dizzy... hang on, processing request...

OpenStudy (anonymous):

And its using the same question as above either use subspace test or explain why its not.

OpenStudy (zugzwang):

Okay. Do you know the subspace test?

OpenStudy (anonymous):

Yup. Test for zero, closure under addition and closure under scalar multiplication right?

OpenStudy (zugzwang):

That's right. Now, is the 0-vector (in this case, the zero polynomial) in W?

OpenStudy (anonymous):

Yh i think.

OpenStudy (zugzwang):

I don't think instructors will accept that :P Let z(x) be the zero polynomial. Then z(x) is in V. Then z(x) = 0, for all x. Then z(2) = 0 = 2(0) = 2[z(1)] There, we showed that the zero polynomial is in W. Care to do the rest of the subspace test? :)

OpenStudy (zugzwang):

Ok, here's the deal. We have to check for closure under addition and closure under scalar multiplication, right? I'll do one of them, you do the other. Which of them do you want me to do?

OpenStudy (anonymous):

you do addition, or do the start and i will try and finish it.

OpenStudy (anonymous):

see here's how i think it works

OpenStudy (zugzwang):

Ok. So, suppose p(x) and q(x) were both in W. Then let r(x) = p(x) + q(x) So that: r(2) = p(2) + q(2) r(1) = p(1) + q(1) Then... r(2) = p(2) + q(2) = 2p(1) + 2q(1) // since p and q are both in W. DON'T FORGET THAT FACT = 2[p(1) + q(1)] // Just "undistributing the 2 = 2r(1) And thus we've shown that r(x) which happens to be the sum of two arbitrary elements of W, is also in W, as it fits the necessary conditions to be in W. Hence, W is closed under vector addition. Now it's your turn :)

OpenStudy (anonymous):

Ok. So suppose c in lR, and u(x) in W. Im stuck. :/ what gets multiplied.

OpenStudy (anonymous):

So suppose c in lR, and u(x) in W. Then let r(x)=cu(x) then cu(2)= c*u(2) = c* 2u(1) = 2cu(1). Thus as r(x) = cu(x), ==> r(2)=2r(1). and thus closure under scalar multiplication.

OpenStudy (anonymous):

Is this correct?

OpenStudy (zugzwang):

Sorry, dozed off, big time o.o. Yep, this is correct